OpenStudy (anonymous):

Evaluate this integral:

OpenStudy (anonymous):

I have tried doing it. I have the answer. Please help with the steps

OpenStudy (experimentx):

let x^2 = u, then 2xdx = du

OpenStudy (anonymous):

Surely it can be done using integration by parts

OpenStudy (anonymous):

why? use a u - sub

OpenStudy (anonymous):

in fact you cannot do it using parts, because you do not know the anti derivative of \(e^{-x^2}\)

OpenStudy (anonymous):

\[u=-x^2,u(1)=-1,u(2)=-4,du=-2xdx,-\frac{1}{2}du=xdx\] \[-\frac{1}{2}\int_{-1}^{-4}udu=\frac{1}{2}\int_{-4}^{-1}udu\] etc